OpenStudy (anonymous):

What is the sum of the interior angles of a dodecagon? Answer 1440° 1620° 1800° 2160°

OpenStudy (anonymous):

150 * 12 = 1800 degrees

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

how did u get 150 and 12

hero (hero):

360/12 = 30 180 - 30 = 150 150*12 = 1800 deg